two coins are tossed once, where
1- E: tail appers on one coin  F: one coin show head
2- E:no tail appears  F: bo head appears
find P(E/F)
 

Grade:12

1 Answers

Arun
5 years ago
For 1st part
 
P(E) = 2/4 = ½
 
P(F) = 2/4 = ½
 
E intersection F = {(h,t),(t,h)}
 
Hence 
P(E intersection F) = 2/4 = ½
 
Now
 
P(E/F) = P(E intersection F)/ P(F) = 1/2 / 1/2 = 1
 
Similarly you can do for 2nd part